If 3x-4,x+4 and 5x+8 are three positive consecutive terms of a gp ..then find the terms

Grade:10

3 Answers

yuron inchit
7 years ago
it is gp so the ratio is same the we can write x-4/3x-4=5x+8/x+4 solve and you`ll get the value of x and put the value in each terms
rusty
7 years ago
As they are the consecutive terms of a gp, we can use the relation b^2=ac. Solving it we will get the value of x and substitute it in the given terms.
gurram meghana
7 years ago
(x+4)2 =(3x-4)(5x+8)
x+8x+16=15x2+24x-20x-32
14x2-4x-48=0
7x2-2x-24=0
7x2-14x+12x-24=0
7x(x-2)+12(x-2)=0
x=2
The required terms are 2,6,18